# Embedding in Recommender Systems: A Survey

### v2

@@ Abstract.
Recommender systems have become an essential component of many online platforms, providing personalized recommendations to users. A crucial aspect is embedding techniques that coverts the high-dimensional discrete features, such as user and item IDs, into low-dimensional continuous vectors and can enhance the recommendation performance. Applying embedding techniques captures complex entity relationships and has spurred substantial research. In this survey, we provide an overview of the recent literature on embedding techniques in recommender systems. This survey covers embedding methods like collaborative filtering, self-supervised learning, and graph-based techniques. Collaborative filtering generates embeddings capturing user-item preferences, excelling in sparse data. Self-supervised methods leverage contrastive or generative learning for various tasks. Graph-based techniques like node2vec exploit complex relationships in network-rich environments. Addressing the scalability challenges inherent to embedding methods, our survey delves into innovative directions within the field of recommendation systems. These directions aim to enhance performance and reduce computational complexity, paving the way for improved recommender systems. Among these innovative approaches, we will introduce Auto Machine Learning (AutoML), hash techniques, and quantization techniques in this survey. We discuss various architectures and techniques and highlight the challenges and future directions in these aspects. This survey aims to provide a comprehensive overview of the state-of-the-art in this rapidly evolving field and serve as a useful resource for researchers and practitioners working in the area of recommender systems. To facilitate the development, evaluation and comparison of embedding-based recommender systems, we provide an open-source algorithm index [^1] https://github.com/Applied-Machine-Learning-Lab/Embedding-in-Recommender-Systems.
doi: XXXXXXX.XXXXXXXccs: Computer science Recommender Systems
@@
Recommender systems have become an integral part of online platforms, providing personalized suggestions to users based on their preferences and behavior (Zangerle and Bauer, 2022). One of the key components of a recommender system is the representation of high-dimensional discrete features in a low-dimensional continuous vector space, known as embeddings. Embeddings have been shown to improve the performance of recommender systems by capturing the complex relationships between entities (e.g. items and users). Although embeddings have been widely employed in recommender systems, there has been a burgeoning body of research dedicated to investigation of them. Various methodologies have been proposed to generate entity embeddings, with the aim of capturing intricate relationships between entities. These methods include collaborative filtering (Rendle, 2010; Abdi et al., 2018; Chen et al., 2022a), self-supervised learning (Yu et al., 2022), and graph-based techniques (Gao et al., 2021). Additionally, Auto Machine Learning (AutoML) (Liu et al., 2020a), hashing techniques (Kang et al., 2020a; Ghaemmaghami et al., 2022), and quantization techniques (Lu and Chang, 2010; Ramanan and Niranjan, 2012) have also been explored as ways to improve the performance and efficiency of embedding-based recommender systems.
@@
Collaborative filtering (CF) is a widely adopted technique in recommender systems (Koren et al., 2021). This method hinges on generating embeddings, which are compact, low-dimensional representations of users and items. These embeddings serve as a means to capture the underlying preferences and intrinsic characteristics of users and items alike. Within collaborative filtering, two prominent strategies for generating embeddings emerge: matrix factorization and factorization machines (Rendle, 2010). The allure of CF lies in its ability to provide accurate recommendations by discerning intricate relationships and interactions among users and items. It notably excels in handling sparse data and effectively tackling the challenge of the cold-start problem, where limited information about new items or users is available (Koren et al., 2021). However, generating embeddings through collaborative filtering (CF) requires extensive training on large datasets, potentially imposing a significant computational burden. Moreover, it is worth noting that CF approaches might exhibit suboptimal performance when dealing with data that is inherently limited or predisposed to biases.
@@
Graph-based methods, such as node2vec (Grover and Leskovec, 2016) and graph autoencoders (Simonovsky and Komodakis, 2018; Ding et al., 2021), harness the rich structural intricacies inherent in item-item and user-user relationships to create embeddings that capture these connections. By adeptly capturing the multifaceted network patterns and associations present in recommendation scenarios, these methods yield embeddings that elegantly encapsulate the existing intricate relationships between entities. These techniques assume particular significance in the settings where there exist an intricate underlying graph structure, as witnessed in social network environments. Furthermore, they furnish a conduit for incorporating auxiliary information through item-item and user-user relationships (Fan et al., 2019b; Chen et al., 2019).
@@
Overall, this survey seeks to offer a thorough panoramic view of embedding applications for recommender systems. We will delve into three categories of widely utilized techniques for embedding generation, encompassing (1) collaborative filtering, (2) self-supervised learning, and (3) graph-based approaches. Furthermore, we will scrutinize integrating (a) AutoML, (b) hashing techniques, and (c) quantization methodologies into the embedding framework. Moreover, in each chapter, we demonstrate existing tools and unveil prospective pathways for future advancement. Finally, we will briefly showcase how embeddings are practically applied in different real-world recommender systems, providing a practical outlook on the methods discussed. To the best of our knowledge, this paper represents the first systematic review of embedding techniques in recommendation systems.
1. Collaborative Filtering Embedding
In this section, we will introduce the strategy of Collaborative Filtering (CF), which is a conventional embedding method in recommender systems. Generally, CF is to learn the similarity between users and items as well as user behavior, then predict potential preferred items for recommendation. There are two typical methods to implement embeddings in CF, Matrix Factorization (MF) scheme and the Factorization Machine (FM) scheme. MF is to factorize the user rating matrix into generalized product of the user feature matrix and item feature matrix, then achieve recommendation according to the preference of users to items, which can be calculated by corresponding user embeddings vectors and item embeddings vectors. However, for simple MF methods, the sparsity of the rating matrix is always a hard issue due to the large amount of one-hot encoded categorical features, and auxiliary features about users and items can not be adopted. To address these issues, FM (Rendle, 2010) is proposed as the initial learning strategy that learns the prediction performance with respect to the similarity of feature embeddings between different features, based on the feature embedding transformed from one-hot embedding.
@@ 1.1. Matrix Factorization Scheme
The singular value decomposition (SVD) is the most common kind of method for matrix factorization, for which the SVD is introduced as a pseudo form in FunkSVD model (Funk, 2006). Intuitively, it involves an approximate factorization of the original rating matrix $\mathbf{R}\in\mathbb{R}^{{m\times n}}$ ($m$ users and $n$ items) to obtain smaller user matrix $\mathbf{U}\in\mathbb{R}^{{m\times d}}$ and item matrix $\mathbf{V}\in\mathbb{R}^{{n\times d}}$, with a hidden embedding dimension $d$:
@@
$\mathbf{U}$ and $\mathbf{V}$ represent unexplainable latent features of the users and the items respectively. The final optimization object is designed as
@@
Since then, many methods with improvements and variations based on FunkSVD (Funk, 2006) were gradually presented. For user factor reduction, NSVD (Paterek, 2007) combines the item factors in a linear way to represent users. SVD++ (Koren, 2008) then replaces representing step with feeding latent user factors to optimize NSVD (Paterek, 2007) by rating through a sparse weight matrix. Then SVDfeature (Chen et al., 2012) was presented as a toolkit that combines rank-model. DELF (Cheng et al., 2018) implements the attention (Vaswani et al., 2017) to automatically differentiate interacted items whose importance is different. BiasSVD (Koren et al., 2009), adds a bias part: user-independent or item-independent factors, to FunkSVD (Funk, 2006). TimeSVD (Koren, 2009) adds time weights by learning a parameter in each period and training with related data.
Beyond direct modification of the FunkSVD framework, several frameworks have also been proposed through SLIM (Ning and Karypis, 2011). SLIM (Ning and Karypis, 2011) introduces a user behavior matrix whose values $r_{ui}$ are binary: 1 for rated item $i$ by user $u$, 0 otherwise. Then replace one of the factorized small matrices. But SLIM (Ning and Karypis, 2011) lacks the ability to compare the similarity of either two items rated by one user at the same time. FISM (Kabbur et al., 2013) solves this by learning the low-dimensional latent vector space without the detailed rating information. But it considers that all the items rated by a user are equal for further prediction. NAIS (He et al., 2018) uses attention to select items with more important interactions automatically, which fixes this issue well. SGNS (Levy and Goldberg, 2014) implements skip-gram (Mikolov et al., 2013c) and negative sampling techniques to factorize a shifted PMI matrix. ConvMF (Kim et al., 2016) integrates the CNN into PMF (Mnih and Salakhutdinov, 2007) and captures contextual information about the document to improve the performance of rating prediction.
@@ 1.2. Factorization Machines
where computation time is linear due to the concept of introducing high-order interactions based on linear regression. The dot product of $k$-dimensional vectors $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ is computed as:
(4) $\mathbf{v}_{i},\mathbf{v}_{j}\right>=\sum^{k}_{f=1}\mathbf{V}_{i,f}\mathbf{V}_{j,f}.$
@@
Parameters $\mathbf{w}=(w_{1},\cdots,w_{n})^{T}\in\mathbb{R}^{n}$ and $\mathbf{V}=(\mathbf{v}_{1},\cdots,\mathbf{v}_{k})^{T}\in\mathbb{R}^{n\times k}$ are used, where $w_{0}$ and $w_{i}$ represent bias and parameter for the $i$-th variable respectively. $k$ signifies the dimension of learned latent embeddings, while $n$ denotes the dimension of the original feature vector. FM facilitates model resolution by replacing $w_{i,j}$ with the dot product of vectors $v_{i}$ and $v_{j}$, effectively achieving the factorization process. When FM exclusively uses user and item IDs with binary interaction outputs, its prediction process aligns with FunkSVD (Zhang, 2022). Nonetheless, FM extends its reach by incorporating the embedding mechanism’s features and extending to diverse feature sets, bolstering flexibility and applicability.
@@
Deep Neural Networks (DNNs) excel in capturing high-order feature interactions (Rendle, 2010). Factorization Neural Networks (FNNs) (Zhang et al., 2016a) enhance this capability by integrating DNNs as the lower layer, enabling the learning of complex interactions that FM’s linear approach lacks. It reduces the need for feature engineering, enhances model expressiveness, and offers inspiration for FM-MLP combinations. Product-based Neural Networks (PNN) (Qu et al., 2016) introduces a product layer comprising linear and nonlinear segments to explicitly calculate the product expression of 2-dimensional feature interactions, while it may perform well with MLPs like FNN (Zhang et al., 2016a).
Neural Factorization Machines (NeuFM) (He and Chua, 2017) follow a similar linear and nonlinear composition, fusing 2-dimensional linear features (derived from a general matrix factorization layer) with high-dimensional nonlinear features from DNNs. Attentional Factorization Machines (AFM) (Xiao et al., 2017) introduce an attention mechanism, learning attention score for feature interaction importance through a neural network. This mechanism reflects varying feature importance, enhancing model sensitivity.
@@
Contrasting the aforementioned method, a fusion of factorization machines (FM) and deep neural networks (DNNs) provides an avenue for alleviating the manual feature engineering burden inherent in certain DNNs. Notably, Google’s Wide & Deep (Cheng et al., 2016) represents a significant advancement in recommendation systems. This model effectively tackles the challenges of memorization and generalization by capturing interactions between low-dimensional and high-dimensional features. However, it still requires artificial feature engineering for the Wide component’s input. Huawei’s DeepFM (Guo et al., 2017) builds upon the Wide & Deep framework, replacing the Wide component with FM. By sharing the embedding part between FM and Deep components, training efficiency improves. Importantly, DeepFM dispenses with pre-training FM and artificial feature engineering, achieving simultaneous learning of low-dimensional and high-dimensional feature interactions.
Further progress, building upon Wide & Deep (Cheng et al., 2016) and DeepFM (Guo et al., 2017), centers on enhancing explicit higher-dimensional feature interactions, particularly through cross-product computations. Deep & Cross Network (DCN) (Wang et al., 2017), an initial approach in this direction, adopts a cross-network instead of FM, replacing the Wide component. While promising, DCN might not universally suit all scenarios. xDeepFM (Lian et al., 2018) achieves simplification by presenting a compressed interaction network (CIN) based on DCN (Wang et al., 2017), and the way of learning feature interaction is changed to a vector-wise idea referred to FM (Rendle, 2010), which means calculating dot-product and then learning the explicit interactions. AutoInt (Song et al., 2019) shares similarities with models like DCN (Wang et al., 2017) and xDeepFM (Lian et al., 2018), all aimed at capturing explicit high-dimensional feature interactions. However, AutoInt differentiates itself by sidestepping the combination of a cross-network and deep neural network. Instead, it directly integrates an interacting layer after the embedding phase.
1.3. Survey and Tool
Survey: Abdi et al.(Abdi et al., 2018) focuses on the field of context-aware recommendation based on matrix factorization, while Chen et al. (Chen et al., 2022a) analyzes and concludes the theory of nonnegative matrix factorization with DNN.

### v3

@@ Abstract.
Recommender systems have become an essential component of many online platforms, providing personalized recommendations to users. A crucial aspect is embedding techniques that convert the high-dimensional discrete features, such as user and item IDs, into low-dimensional continuous vectors, which can enhance the recommendation performance. Embedding techniques have revolutionized the capture of complex entity relationships, generating significant research interest. This survey presents a comprehensive analysis of recent advances in recommender system embedding techniques. We examine centralized embedding approaches across matrix, sequential, and graph structures. In matrix-based scenarios, collaborative filtering generates embeddings that effectively model user-item preferences, particularly in sparse data environments. For sequential data, we explore various approaches including recurrent neural networks and self-supervised methods such as contrastive and generative learning. In graph-structured contexts, we analyze techniques like node2vec that leverage network relationships, along with applicable self-supervised methods. Our survey addresses critical scalability challenges in embedding methods and explores innovative directions in recommender systems. We introduce emerging approaches, including AutoML, hashing techniques, and quantization methods, to enhance performance while reducing computational complexity. Additionally, we examine the promising role of Large Language Models (LLMs) in embedding enhancement. Through detailed discussion of various architectures and methodologies, this survey aims to provide a thorough overview of state-of-the-art embedding techniques in recommender systems, while highlighting key challenges and future research directions. To facilitate development, evaluation, and comparison of embedding-based recommender systems, we provide an open-source repository  [^1] https://github.com/Applied-Machine-Learning-Lab/Embedding-in-Recommender-Systems.
Embedding Learning, Recommender Systems, Graph Neural Networks, Matrix Factorization, Sequential Recommendation, Self-supervised Learning, AutoML, Hash Embedding, Quantization, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systemsccs: Networksccs: Applied computing
1. Introduction
@@
Recommender systems have become an integral part of online platforms, providing personalized suggestions to users based on their preferences and behavior (Zangerle and Bauer, 2022; Wang et al., 2025c, b). One of the key components of a recommender system is the representation of high-dimensional discrete features in a low-dimensional continuous vector space, known as embeddings. Embeddings have been shown to improve the performance of recommender systems by capturing the complex relationships between entities (e.g., items and users). Although embeddings have been widely employed in recommender systems, there has been a burgeoning body of research dedicated to the investigation of them. Various methodologies have been proposed to generate entity embeddings, with the aim of capturing intricate relationships between entities. These methods include collaborative filtering (Rendle, 2010; Abdi et al., 2018; Chen et al., 2022a), self-supervised learning (Yu et al., 2022), and graph-based techniques (Gao et al., 2021). Additionally, Auto Machine Learning (AutoML) (Liu et al., 2020a), hashing techniques (Kang et al., 2020a; Ghaemmaghami et al., 2022), and quantization techniques (Lu and Chang, 2010; Ramanan and Niranjan, 2012) have also been explored as ways to improve the performance and efficiency of embedding-based recommender systems.
@@
Collaborative filtering (CF) is a widely adopted technique in recommender systems (Koren et al., 2021) for matrix data format. This method hinges on generating embeddings, which are compact, low-dimensional representations of users and items. These embeddings serve as a means to capture the underlying preferences and intrinsic characteristics of users and items alike. Within collaborative filtering, two prominent strategies for generating embeddings emerge: matrix factorization and factorization machines (Rendle, 2010). The allure of CF lies in its ability to provide accurate recommendations by discerning intricate relationships and interactions among users and items. It notably excels in handling sparse data and effectively tackling the challenge of the cold-start problem, where limited information about new items or users is available (Koren et al., 2021). However, generating embeddings through collaborative filtering (CF) requires extensive training on large datasets, potentially imposing a significant computational burden. Moreover, it is worth noting that CF approaches might exhibit suboptimal performance when dealing with data that is inherently limited or predisposed to biases.
@@
Structure-based methods, such as node2vec (Grover and Leskovec, 2016) and graph autoencoders (Simonovsky and Komodakis, 2018; Ding et al., 2021) in graph modeling, and sequence modeling techniques like LSTM, GRU, and Transformer, harness rich structural intricacies to create embeddings that capture inherent connections. By effectively capturing the multifaceted network patterns and associations present in recommendation scenarios, these methods yield embeddings that encapsulate the intricate relationships between entities. These techniques are particularly significant in settings with complex underlying structures, such as social network environments and sequential recommendation scenarios like e-commerce browsing history and video watching sequences. Furthermore, they provide a means to incorporate auxiliary information through item-item and user-user relationships (Fan et al., 2019).
@@
Overall, this survey seeks to offer a thorough panoramic view of embedding applications for recommender systems. We will delve into three categories of widely utilized techniques for embedding generation, encompassing (1) collaborative filtering, (2) self-supervised learning, and (3) graph-based approaches. Furthermore, we will scrutinize integrating (a) AutoML, (b) hashing techniques, and (c) quantization methodologies into the embedding framework. Moreover, in each chapter, we demonstrate existing surveys and unveil prospective pathways for future advancement. Finally, we will briefly showcase how embeddings are practically applied in Click-Through-Rate (CTR) predictions and Large Language Models (LLMs) based recommendations, providing a practical outlook on the embedding usage. To the best of our knowledge, this paper represents the first systematic review of embedding techniques in recommendation systems.
2. Embedding Learning of Matrix Format
In this section, we will explore traditional embedding methods in recommender systems, specifically Collaborative Filtering (CF) (Sarwar et al., 2001). CF primarily focuses on learning the similarities between users and items as well as user behavior to predict potential item preferences for recommendations, commonly employing matrix format modeling. The two typical methods used to implement embeddings in CF are Matrix Factorization (MF) and Factorization Machines (FM). Matrix Factorization (Koren, 2008) involves decomposing the user rating matrix into a generalized product of the user feature matrix and the item feature matrix, thereby facilitating recommendations based on user preferences for items. These preferences are calculated using the corresponding user and item embedding vectors. However, simple MF methods face challenges with the sparsity of the rating matrix due to the extensive use of one-hot encoded categorical features, and they cannot utilize auxiliary features about users and items. To address these issues, FM (Rendle, 2010) is introduced as an initial learning strategy that enhances prediction performance by learning the similarities between feature embeddings derived from one-hot encoding. By emphasizing matrix formats, this method not only highlights its mathematical structure but also its pivotal role in embedding learning, especially in handling large-scale sparse data. A summary of the methods discussed in this section is presented in Table 1. Generally, CF aims to learn the similarity between users and items as well as user behavior, then predict potential preferred items for recommendation.
Then, MF implements machine learning models to fill the missing ratings of the original user-item matrix for the final recommendation task.
@@ 2.1. Matrix Factorization Scheme
Table 1. Embedding Learning of Matrix Format
Embedding Learning of Matrix Format Matrix Factorization Scheme (§2.1) SVD-based Enhancements: FunkSVD (Funk, 2006), NSVD (Paterek, 2007), SVD++ (Koren, 2008), SVDfeature (Chen et al., 2012), DELF (Cheng et al., 2018), BiasSVD (Koren et al., 2009), TimeSVD (Koren, 2009)
Item Similarity Models: SLIM (Ning and Karypis, 2011), FISM (Kabbur et al., 2013), NAIS (He et al., 2018), SGNS (Levy and Goldberg, 2014)
Content-aware Models: ConvMF (Kim et al., 2016), ConvSeq-MF (Khan et al., 2025)
Factorization Machines (§2.2) Foundation Model: Factorization Machines (FM) (Rendle, 2010)
Deep Learning Hybrids: FNN (Zhang et al., 2016), PNN (Qu et al., 2016), NeuFM (He and Chua, 2017), AFM (Xiao et al., 2017)
Wide & Deep Architectures: Wide & Deep (Cheng et al., 2016), DeepFM (Guo et al., 2017), DCN (Wang et al., 2017), xDeepFM (Lian et al., 2018)
MF aims to connect, represent the users and items according to the latent feature factorized. The singular value decomposition (SVD) is the most common kind of method for matrix factorization, for which the SVD is introduced as a pseudo form in FunkSVD model (Funk, 2006). Intuitively, it involves an approximate factorization of the original rating matrix $\mathbf{R}\in\mathbb{R}^{{m\times n}}$ ($m$ users and $n$ items) to obtain smaller user matrix $\mathbf{U}\in\mathbb{R}^{{m\times d}}$ and item matrix $\mathbf{V}\in\mathbb{R}^{{n\times d}}$, with a hidden embedding dimension $d$:
@@
$\mathbf{U}$ and $\mathbf{V}$ represent unexplainable latent features of the users and the items, respectively. The final optimization object is
@@
Since then, many methods with improvements and variations based on FunkSVD (Funk, 2006) have been gradually presented. For user factor reduction, NSVD (Paterek, 2007) combines the item factors in a linear way to represent users. SVD++ (Koren, 2008) then replaces the representing step with feeding latent user factors to optimize NSVD (Paterek, 2007) by rating through a sparse weight matrix. Then, SVDfeature (Chen et al., 2012) was presented as a toolkit that combines a rank model. DELF (Cheng et al., 2018) implements the attention (Vaswani et al., 2017) to automatically differentiate interacted items whose importance is different. BiasSVD (Koren et al., 2009) adds a bias part: user-independent or item-independent factors, to FunkSVD (Funk, 2006). TimeSVD (Koren, 2009) adds time weights by learning a parameter in each period and training with related data. While these traditional matrix factorization variants have established strong foundations, recent research has identified new challenges and opportunities in practical deployment. Modern applications demand not only accuracy but also security, robustness, and enhanced representation learning capabilities (Nguyen et al., 2024; Wu et al., 2025). Addressing these emerging needs, Shams et al. (Shams and Leith, 2025) tackle the critical issue of security in production recommender systems, introducing attack detection mechanisms that monitor item vector shifts to identify malicious manipulations. This work highlights the gap between theoretical matrix factorization models and real-world deployment requirements.
Beyond direct modification of the FunkSVD framework, several frameworks have also been proposed through SLIM (Ning and Karypis, 2011). SLIM (Ning and Karypis, 2011) introduces a user behavior matrix whose values $r_{ui}$ are binary: 1 for rated item $i$ by user $u$, 0 otherwise. Then replace one of the factorized small matrices. But SLIM (Ning and Karypis, 2011) lacks the ability to compare the similarity of either Two items were rated by one user at the same time. FISM (Kabbur et al., 2013) solves this by learning the low-dimensional latent vector space without the detailed rating information. But it considers that all the items rated by a user are equal for further prediction. NAIS (He et al., 2018) uses attention to select items with more important interactions automatically, which fixes this issue well. SGNS (Levy and Goldberg, 2014) implements skip-gram (Mikolov et al., 2013b) and negative sampling techniques to factorize a shifted PMI matrix. ConvMF (Kim et al., 2016) integrates the CNN into PMF (Mnih and Salakhutdinov, 2007) and captures contextual information about the document to improve the performance of rating prediction. Building upon ConvMF’s integration of textual content, the evolution toward comprehensive multimodal matrix factorization continues to gain momentum. Khan et al. (Khan et al., 2025) advance this direction with ConvSeq-MF, which systematically combines convolutional neural networks with sequence modeling while leveraging both user and item auxiliary content information. Furthermore, the integration of large language models into matrix factorization has opened new possibilities for semantic understanding. Liu et al. (Liu et al., 2025a) demonstrate how LLMs can enhance traditional collaborative filtering by providing semantic aspect-aware review exploitation, bridging the gap between textual content and latent factor models.
@@ 2.2. Factorization Machines
where computation time is linear due to the concept of introducing high-order interactions based on linear regression. The dot product of $k$-dimensional vectors $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ is:
(4) $\left<\mathbf{v}_{i},\mathbf{v}_{j}\right>=\sum^{k}_{f=1}\mathbf{V}_{i,f}\mathbf{V}_{j,f}.$
@@
Parameters $\mathbf{w}=(w_{1},\cdots,w_{n})^{T}\in\mathbb{R}^{n}$ and $\mathbf{V}=(\mathbf{v}_{1},\cdots,\mathbf{v}_{k})^{T}\in\mathbb{R}^{n\times k}$ are used, where $w_{0}$ and $w_{i}$ represent bias and parameter for the $i$-th variable respectively. $k$ signifies the dimension of learned latent embeddings, while $n$ denotes the dimension of the original feature vector. FM facilitates model resolution by replacing $w_{i,j}$ with the dot product of vectors $v_{i}$ and $v_{j}$, effectively achieving the factorization process. When FM exclusively uses user and item IDs with binary interaction outputs, its prediction process aligns with FunkSVD (Zhang, 2022). Nonetheless, FM extends its reach by incorporating the embedding mechanism’s features and extending to diverse feature sets, bolstering flexibility and applicability. While factorization machines effectively handle sparse features through second-order interactions, they face computational challenges when dealing with numerous categorical fields in real-time scenarios. Recent work has explored more efficient approaches for handling field interactions in latency-critical applications. Low-rank field-weighted factorization machines (Shtoff et al., 2024) explore the low-rank decomposition of field interaction matrices to reduce computational complexity while maintaining modeling expressiveness for online serving scenarios. Moreover, the limitation of modeling only pairwise user-item interactions has been highlighted, as real-world systems often exhibit complex, higher-order feature dependencies. To address this, GraphFM (Wu et al., 2025) introduces a graph-based factorization machine framework that explicitly models higher-order feature interactions via graph neural networks, significantly extending the expressive power of traditional FM models.
@@
Deep Neural Networks (DNNs) excel in capturing high-order feature interactions (Rendle, 2010). Factorization Neural Networks (FNNs) (Zhang et al., 2016) enhance this capability by integrating DNNs as the lower layer, enabling the learning of complex interactions that FM’s linear approach lacks. FNN (Zhang et al., 2016) reduces dimensionality by transforming sparse features using supervised learning embedding layers. It reduces the need for feature engineering, enhances model expressiveness, and offers inspiration for FM-MLP combinations. Product-based Neural Networks (PNN) (Qu et al., 2016) introduces a product layer comprising linear and nonlinear segments to explicitly calculate the product expression of 2-dimensional feature interactions, while it may perform well with MLPs like FNN (Zhang et al., 2016). Neural Factorization Machines (NeuFM) (He and Chua, 2017) follow a similar linear and nonlinear composition, fusing 2-dimensional linear features (derived from a general matrix factorization layer) with high-dimensional nonlinear features from DNNs. Attentional Factorization Machines (AFM) (Xiao et al., 2017) introduce an attention mechanism, learning an attention score for feature interaction importance through a neural network. This mechanism reflects varying feature importance, enhancing model sensitivity.
@@
Contrasting the aforementioned method, a fusion of factorization machines (FM) and deep neural networks (DNNs) provides an avenue for alleviating the manual feature engineering burden inherent in certain DNNs. Notably, Google’s Wide & Deep (Cheng et al., 2016) represents a significant advancement in recommendation systems. This model effectively tackles the challenges of memorization and generalization by capturing interactions between low-dimensional and high-dimensional features. However, it still requires artificial feature engineering for the Wide component’s input. Huawei’s DeepFM (Guo et al., 2017) builds upon the Wide & Deep framework, replacing the Wide component with FM. By sharing the embedding part between FM and Deep components, training efficiency improves. Importantly, DeepFM dispenses with pre-training FM and artificial feature engineering, achieving simultaneous learning of low-dimensional and high-dimensional feature interactions. Further progress, building upon Wide & Deep (Cheng et al., 2016) and DeepFM (Guo et al., 2017), centers on enhancing explicit higher-dimensional feature interactions, particularly through cross-product computations. Deep & Cross Network (DCN) (Wang et al., 2017), an initial approach in this direction, adopts a cross-network instead of FM, replacing the Wide component. While promising, DCN might not universally suit all scenarios. xDeepFM (Lian et al., 2018) achieves simplification by presenting a compressed interaction network (CIN) based on DCN (Wang et al., 2017), and the way of learning feature interaction is changed to a vector-wise way. However, AutoInt differentiates itself by sidestepping the combination of a cross-network and a deep neural network. Instead, it directly integrates an interacting layer after the embedding phase.
2.3. Surveys and Future Directions